when two equal resistances are connected in series and parallel combination. find equivalent resistance in both cases.
1
Expert's answer
2020-09-17T14:29:44-0400

Let resistance of each resistor be R.

When these are connected in series then there equivalent is given by "R_s = R + R = 2R"


When these are connected in parallel then equivalent resistance will be "R_p = \frac{RR}{R+R} = \frac{R}{2}"